OpenStudy (anonymous):

PLEASE HELP domain for y=1?

OpenStudy (anonymous):

\(y=1\) is completely independent of \(x\), which means regardless of the value of \(x\), \(y\) will always be 1. So the domain is all numbers.

OpenStudy (jdoe0001):

another way to say it will be \(\begin{matrix} x = \pm2 & y =1\\ x = \pm 2,000 & y =1\\ x = \pm 2,000,000 & y =1\\ x = \pm 2,000,000,000 & y =1\\ x = -\infty & y =1\\ x = +\infty & y =1\\ \end{matrix}\)

OpenStudy (anonymous):

Of course, that's assuming we're dealing with \(y\) as a function of \(x\)...